# Writing – review & editing, Writing – original draft, Visualization, Data curation ,Investigation, Formal analysis, Validation, Software, Methodology, Conceptualization

### v2

@@ Abstract
Machine learning models trained on tabular data are vulnerable to adversarial attacks, even in realistic scenarios where attackers only have access to the model’s outputs. Since tabular data contains complex interdependencies among features, it presents a unique challenge for adversarial samples which must maintain coherence and respect these interdependencies to remain indistinguishable from benign data. Moreover, existing attack evaluation metrics—such as the success rate, perturbation magnitude, and query count—fail to account for this challenge. To address those gaps, we propose a technique for perturbing dependent features while preserving sample coherence. In addition, we introduce Class-Specific Anomaly Detection (CSAD), an effective novel anomaly detection approach, along with concrete metrics for assessing the quality of tabular adversarial attacks. CSAD evaluates adversarial samples relative to their predicted class distribution, rather than a broad benign distribution. It ensures that subtle adversarial perturbations, which may appear coherent in other classes, are correctly identified as anomalies. We integrate SHAP explainability techniques to detect inconsistencies in model decision-making, extending CSAD for SHAP-based anomaly detection. Our evaluation incorporates both anomaly detection rates with SHAP-based assessments to provide a more comprehensive measure of adversarial sample quality. We evaluate various attack strategies, examining black-box query-based and transferability-based gradient attacks across four target models. Experiments on benchmark tabular datasets reveal key differences in the attacker’s risk and effort and attack quality, offering insights into the strengths, limitations, and trade-offs faced by attackers and defenders. Our findings lay the groundwork for future research on adversarial attacks and defense development in the tabular domain.
@@ 1 Introduction
Machine learning (ML) models trained on tabular data are widely deployed in diverse industry sectors including finance (e.g., for computing credit risk) [12, 37], healthcare (e.g., for analyzing administrative claims and patient registry data) [13], and energy (e.g., for predicting energy and electricity consumption) [30, 43]. However, recent research has highlighted their vulnerability to adversarial attacks, in which malicious data samples are used to mislead the model into producing incorrect outputs [2, 8, 36, 21]. For instance, an applicant may slightly increase their reported income (influencing the “income level" feature), crossing a decision threshold and thereby changing the model’s decision from ‘reject’ to ‘approve.’ Such adversarial attacks can even be conducted in realistic black-box settings, where the attacker has no knowledge of the model’s internals and can only query the model to receive outputs.
@@
While extensive research has been performed in the image [50, 52, 26, 42], audio [31], text [50, 35, 20], and graph data  [50, 46] domains, the tabular domain remains relatively underexplored. This paper addresses several key challenges in the realm of adversarial attacks, focusing on tabular data.
Crafting Coherent and Consistent Adversarial Samples: Crafting adversarial samples in the tabular domain presents unique challenges compared to other data types. Unlike domains such as images, audio, and text, which contain homogeneous feature types with well-defined dependencies (spatial in images, temporal in audio, and semantic in text), the tabular domain is characterized by heterogeneous feature types, diverse data constraints, and complex interdependencies. Correlations in tabular data can also be complicated. For instance, in a loan dataset, a high credit score strongly correlates with approval, but for a low credit score, approval depends on other factors like income or debt, making the correlation conditional.
@@
Models trained on tabular data often rely on interdependent features, where the values of some features (dependent features) are influenced by others (influencing features). Therefore, in adversarial attacks, perturbations to dependent features must align with the influencing features to preserve these constraints and interdependencies, maintaining coherence and consistency to avoid detection. For instance, the BMI is derived from height and weight, so when altering these influencing features (such as height or weight) during an attack, it is essential to ensure that the BMI value remains consistent with the changes.
Multiple dependent features further complicate the process of maintaining this alignment. Furthermore, while some dependencies are well-defined (as is the case with the BMI), others may be non-linear, indirect, or influenced by features unknown to the attacker. For instance, the feature “employee performance rating," used in a model predicting company efficiency, depends on factors such as project completion rate, peer feedback, client satisfaction scores, and overtime hours. In practice, the attacker may not have access to all influential features or know exactly how these features contribute to the employee performance rating. These complex, often domain-specific interdependencies and feature correlations make it much harder to maintain consistency and coherence when crafting adversarial samples in the tabular domain.
@@
Evaluation Criteria for Distinguishable Adversarial Samples: Existing evaluation methods for adversarial samples primarily focus on measuring the perturbation magnitude using $L_{p}$ norms, which are intended to indicate the level of detectability (i.e., the risk that the attack will be exposed). However, in the context of tabular data, simply limiting perturbation magnitudes is insufficient. Effective adversarial attacks must ensure that samples remain coherent, consistent, and in-distribution to evade detection. While previous work has acknowledged these challenges (see  Section 2.3), there is a lack of concrete evaluation methods for quantifying the distinguishability of adversarial samples in the tabular domain.
Empirical Analysis of Different Attack Strategies: While adversarial attacks have been extensively studied in other domains, the analysis of black-box attack strategies in the tabular domain has been limited. Query-based and transferability-based are common black-box attack strategies.
@@
In query-based attacks [8], it is assumed that the attacker has unlimited access to the target model’s predictions. By repeatedly querying the model, he attacker is able to iteratively optimize an adversarial example to induce misclassification. Transferability-based attacks [36, 21] operate under a more constrained threat model, where the attacker lacks direct access to the target model’s outputs but is assumed to have access to a large amount of reliable data to train an accurate surrogate model. In this case, the attacker optimizes the adversarial sample to mislead the surrogate model and then uses this sample to mislead the target model with a single query.
@@
Adversarial attacks are used by both defenders and attackers both of whom may choose different attack strategies based on their goals, resources, and constraints. Defenders who are interested in assessing their model’s vulnerability often use adversarial attacks for automated penetration testing or adversarial training, which involves creating adversarial samples to improve the model’s robustness [11, 23]. Since defenders have full access to their model and can make unlimited queries, query-based attacks are a practical option. In contrast, attackers aiming to minimize the number of queries performed and avoid detection, might prefer transferability-based attacks that use gradient methods on neural networks, as they capture underlying feature interactions necessary for subtle perturbations. Understanding how and whether query-based attacks reflect other types of attacks, such as transferability-based gradient attacks, is crucial when using them to improve model defenses.
@@
In this paper, we aim to advance adversarial research in the tabular domain by addressing existing gaps in the literature. We compare the effectiveness of various attack strategies against classification models to provide insights for both attackers and defenders. Additionally, we developed a novel technique for consistent perturbation of dependent features and concrete evaluation criteria for assessing attack quality. We perform a detailed examination of both black-box query- and transferability-based gradient attacks, tailored to respect tabular distribution constraints.
Our proposed technique for perturbing dependent features aims to address the challenge of crafting coherent and consistent adversarial samples. This technique uses regression models to correct dependent features during the attack process, ensuring that they maintain realistic relationships with interdependent features and remain consistent with the values of other features (see Section 3).
To assess the distinguishability and quality of adversarial samples (see Section 3), we introduce Class-Specific Anomaly Detection (CSAD), an approach that measures the anomaly detection rate of adversarial samples relative to the distribution of benign samples from the specific class predicted by the target model. CSAD leverage existing anomaly detection methods to detect anomalies when inconsistencies are present within class-specific distributions. Additionally, we propose complementary criteria that use explainability-based analysis to examine how adversarial samples influence the model’s decision-making process. This analysis uses SHapley Additive exPlanations (SHAP)[33], a widely adopted explainability (XAI) technique[3], to quantify the inconsistencies induced by adversarial perturbations in the model’s internal reasoning.
Our comparison of attack strategies’ performance factors integrates both traditional metrics of the attacker’s risk (success rate, query count, and $L_{p}$ norms) and effort and the new evaluation criteria introduced in this paper (sample anomaly detection and feature importance consistency), providing a thorough assessment of adversarial sample quality and attack effectiveness.
Four different tree-based target models are used to explore and evaluate the attack strategies (process illustrated in  Fig. 1) on three datasets: the Hateful Users on Twitter [44] and Intensive Care Unit [22] public datasets, which are benchmark datasets in the tabular domain, and the Video Transmission Quality (VideoTQ) dataset, which is a proprietary dataset. Seven attacks are examined in our evaluation: two well-known query-based attacks (the boundary and HopSkipJump attacks [6, 9] that can be adapted for tabular data [8]) and five transferability-based gradient attacks that varied in terms of their feature selection techniques (a basic attack [36] in which features were randomly selected and four attacks in which features were selected based on their importance [21]).
This work sheds light on the key trade-offs between these attacks for the tabular domain, which can guide the selection of attack strategies and motivate the development of new defense approaches tailored to the unique properties of this domain.
The contributions of our work can be summarized as follows:
@@
  • We propose a novel technique for perturbing interdependent features in tabular data, leveraging regression models to preserve coherence and ensure feature consistency, overcoming a key challenge in tabular adversarial attacks.
@@
  • We introduce CSAD, an approach that assesses anomalies relative to class-specific distributions rather than global thresholds, which improve sensitivity to inconsistencies among features.
@@
  • We developed comprehensive dual evaluation criteria based on anomaly detection assessment of the sample values and feature importance, providing a core principled framework to quantify both the quality of tabular adversarial examples.and their impact on model decision-making.
@@
  • To the best of our knowledge, this study is the first to systematically compare adversarial attack strategies’ performance factors in the tabular domain, providing critical insights into their effectiveness, limitations, and corresponding defense strategies.
  • The resources of this study (preprocessed data, code, models, and adversarial samples) will be made publicly available to promote future research and reproducibility.
@@
The remainder of this paper is organized as follows: In Section 2, we discuss related work. Section 3 introduces our methodological framework, followed by Section 4 and Section 5, which present the experimental setup and provide a thorough analysis of the examined query- and transferability-based attack behavior. Finally, Section 6 and Section 7 conclude with insights for attackers and defenders within tabular data domains.
@@ 2 Background and Related Work
2.1 State-of-the-Art Adversarial Attacks in Tabular Data Domains
@@
Szegedy et al. [47] introduced the concept of an adversarial attack in which an attacker provides a malicious data sample aimed at misleading the ML model and causing it to produce incorrect output. Their adversarial sample was crafted by performing subtle changes to an image, which were imperceptible to the human eye but misled the target model. However, research conducted in tabular subdomains, such as fraud detection and recommendation systems [2, 8, 14], has highlighted additional critical challenges within the tabular domain. These challenges include the constraints of editable and non-editable features (e.g., ID, date of birth), data imbalance, and the presence of noncontinuous features (e.g., categorical features). The authors of these studies also mentioned the complexity of perturbing discrete values while maintaining the semantic integrity of the sample, as well as capturing nuances when designing attacks. This discussion was further elaborated on in [36].
Most recent adversarial attacks were demonstrated under black-box conditions (i.e., the attacker has no prior knowledge of the target model and only has the ability to query it [34]). The main types of black-box attacks are query-based attacks and transferability-based attacks.
In query-based attacks, the attacker assumes that the target model can be queried multiple times, and when queried, the attacker receives the model’s confidence or classification. The attacker iteratively adjusts the sample’s values, based on the model’s output, until the optimization objective is fulfilled - the target model is misled and produces the desired output. Decision-based attacks are examples of black-box query-based attacks [7]. These attacks typically involve making a large initial modification to the original sample, such that the target model produces an incorrect prediction. Then the attacker optimizes the modified sample’s values so they are as close as possible to the original sample while still misleading the target model. For evaluation, we implemented two decision-based attacks: The boundary attack and the HopSkipJump attack, which differ in terms of the extent of the initial changes made to the sample and the optimization process performed [6, 9]. Modifications to these decision-based attacks for incorporating tabular data validity and editability constraints were made by [8].
In transferability-based attacks, the attacker trains a surrogate model, crafts an adversarial sample using a white-box attack on the surrogate model, and then utilizes the adversarial sample crafted to attack the target model in a single query [39].
An untargeted, transferability-based gradient attack designed to address tabular constraints was introduced in previous work [36, 21]. This architecture is based on NN and incorporates an embedding function to preserve feature correlations and ensure value consistency by minimizing $L_{2}$ distance and adjusting correlated features. The attack process iteratively perturbs features to reduce $L_{0}$ (i.e., the number of modified features) while applying validity and editability constraints, demonstrating its applicability in real-world scenarios. Another work [25] introduced a transferability-based attack that considers feature distribution in the perturbation process, addressing the tabular distribution challenge while excluding validity constraints.
2.2 Existing Methods for Evaluating Adversarial Attacks
Commonly used metrics for assessing the performance of adversarial attacks include:
  1. 1.
    Attacker’s risk. The attacker’s risk is assessed by the attack success rate [19, 2], the number of queries made to the target model [24, 48], and the extent to which the adversarial sample has been distorted, consequently becoming distinguishable. The distortion is typically measured using $L_{p}$ norms such as $L_{0}$, $L_{2}$, and $L_{\infty}$, which quantify the magnitude of perturbations applied to the original samples [19, 29, 15, 2, 36, 21], and can also be evaluated by computing the average error between the original sample and the corresponding adversarial samples[18].
  2. 2.
    Attacker’s effort. The attacker’s effort is assessed by the time required to create the adversarial sample, as well as the amount of data and time needed to train the surrogate model [1].
2.3 Indistinguishability in the Tabular Domain
While most metrics used to assess the attacker’s effort and risk are consistent across various domains including tabular data, the detectability dimension of adversarial samples should be measured differently for tabular data domains. In the computer vision domain, for example, an adversarial sample is considered to be indistinguishable if the human eye is unable to recognize the difference between the adversarial sample and the original sample [47, 19, 40, 29]. It is commonly assessed by measuring the distance between the original and perturbed images using $L_{2}$ norm.
However, in the tabular domain, “human eye" inspection is less meaningful, as only domain experts can accurately examine such differences [2, 8]. Furthermore, when looking at tabular data samples, a domain expert would “select" a limited number of features to examine and determine whether the sample is real or manipulated by an adversary [2] based on those features. Given the above, the authors of [2, 8] proposed evaluating the distinguishability of a tabular adversarial sample based on the number of “important" features (i.e., features that a domain expert would likely examine) altered in the adversarial attack. Accordingly, an adversarial sample is considered more perceptible as the number of changes in important features increases. However, this evaluation method is neither objective nor practical, as it requires a domain expert to manually assess each sample individually. Moreover, it does not consider the fact that altering even a small number of features can cause inconsistency with the rest of the sample, making it distinguishable from the benign data distribution and detectable by anomaly-based methods.
Prior research highlighted the significance of both consistency [36, 25] and validity [36, 18] when analyzing the quality of adversarial samples. Validity can be assessed by examining whether each feature value falls within a valid range [2, 36] and by measuring the sample’s deviation from the distribution of benign data [36]. However, no concrete metric has been used to quantify the coherence and consistency of adversarial samples, which are influenced by complex correlations between feature values [36].
We address this gap by proposing two criteria that assess the distinguishability of adversarial samples from benign ones using an “anomaly detection rate," which provides an indirect evaluation of their coherence and consistency. We assume that adversarial samples lacking structural coherence or consistency with the benign data distribution are more likely to exhibit anomalies, making them both distinguishable and detectable.

### v3

@@ Abstract
Machine learning models trained on tabular data are vulnerable to adversarial attacks, even in realistic scenarios where attackers only have access to the model’s outputs. Since tabular data contains complex dependencies among features, it presents a unique challenge for adversarial samples which must maintain coherence and respect these dependencies to remain indistinguishable from benign data. Moreover, existing attack evaluation metrics—such as the success rate, perturbation magnitude, and query count-fail to account for this challenge. To address these gaps, we propose a technique for perturbing dependent features while preserving sample coherence. In addition, we introduce Class-Specific Anomaly Detection (CSAD), an effective novel anomaly detection approach, along with concrete metrics for assessing the quality of tabular adversarial attacks. CSAD evaluates adversarial samples relative to their predicted class distribution, rather than a broad benign distribution. This ensures that subtle adversarial perturbations, which may appear coherent in other classes, are correctly identified as anomalies. We extend CSAD for importance-based anomaly detection by integrating SHAP explainability techniques to detect inconsistencies in model decision-making. Our evaluation of adversarial sample quality incorporates both anomaly detection rates and importance-based assessments to provide a more comprehensive measure. We evaluate various attack strategies, examining black-box query-based and transferability-based gradient attacks across four target models. Experiments on benchmark tabular datasets reveal key differences in the attacker’s risk and effort and attack quality, offering insights into the strengths, limitations, and trade-offs faced by attackers and defenders. Our findings lay the groundwork for future research on adversarial attacks and defense development in the tabular domain.
@@ 1 Introduction
Machine learning (ML) models trained on tabular data are widely deployed in diverse industry sectors where they are used for a wide range of tasks, including credit risk assessment in finance [1, 2], the analysis of administrative claims and patient registries in healthcare [3], and the prediction of energy consumption patterns in energy [4, 5]. However, recent research has highlighted their vulnerability to adversarial attacks, in which malicious data samples are used to mislead the model into producing incorrect outputs [6, 7, 8, 9]. Such attacks can even be conducted in realistic black-box settings, where the attacker has no knowledge of the model’s internals and can only query the model to receive outputs.
The core challenge in adversarial attacks on tabular data lies in maintaining realistic data relationships while achieving misclassification [8, 9]. Unlike domains such as images, audio, or text, which contain homogeneous feature types with well-defined dependencies (spatial, temporal, or semantic respectively), the tabular domain presents unique complexities. Features in tabular datasets are heterogeneous, subject to diverse constraints, and exhibit complex interdependencies that must be preserved to avoid detection [7, 8, 9].
Consider a loan approval scenario: an applicant may slightly increase their reported income to cross a decision threshold, changing the model’s decision from ‘reject’ to ‘approve.’ However, this modification must remain consistent with other financial indicators to appear legitimate. This illustrates the fundamental challenge—crafting adversarial samples that are both effective and undetectable.
@@
While extensive research has been performed on images [10, 11, 12, 13], audio [14], text [10, 15, 16], and graph data  [10, 17], the tabular domain remains relatively underexplored. This paper addresses three critical gaps in tabular adversarial research.
First, research on preserving feature dependencies and coherence in adversarial attacks is limited.
@@
Models trained on tabular data often rely on interdependent features, where the values of some features (dependent features) are influenced by others (influencing features). Therefore, in adversarial attacks, perturbations to dependent features must align with the influencing features to preserve these constraints and interdependencies, maintaining coherence and consistency to avoid detection. For instance, the body mass index (BMI) is derived from height and weight [18], so when altering one of these influencing features (i.e., height or weight) during an attack, it is essential to ensure that the BMI value remains mathematically consistent with the changes.
In practice, correlations in tabular data can be complex. In the loan approval scenario, for example, a high credit score strongly correlates with approval, but for a low credit score, approval depends on other factors like income or debt, making the correlation conditional. In another practical case, such as employee performance ratings which depend on multiple factors such as project completion rates, client satisfaction scores, and peer feedback, the dependencies are more complex, and may be non-linear, indirect, or influenced by features unknown to the attacker. Preserving such dependencies during perturbation poses a significant challenge.
@@
Second, current attack evaluation methods are not adapted to tabular data. Existing approaches primarily focus on measuring the perturbation magnitude using $L_{p}$ norms [6, 7, 8], which aim to indicate the detectability risk (i.e., the risk that the attack will be exposed). However, in the context of tabular data, limiting perturbation magnitudes alone is insufficient. Effective adversarial attacks must ensure that samples remain coherent, consistent, and in-distribution to evade detection. While previous work has acknowledged these challenges (see  Section˜2.2), there is a lack of concrete evaluation methods for quantifying the distinguishability of adversarial samples in the tabular domain.
@@
Third, empirical analysis of attack strategies in the tabular domain is limited. Two common black-box attack strategies exist: query-based attacks, where attackers are assumed to have unlimited access to model predictions and iteratively optimize the adversarial samples by repeatedly querying the model [7], and transferability-based attacks, where attackers lack direct access to the target model’s outputs but train surrogate models using available data [8, 9]. In the latter case, the attacker optimizes the adversarial sample to mislead the surrogate model and then uses this sample to mislead the target model with a single query.
@@
Both attackers and defenders need to understand the practical trade-offs between these approaches, yet systematic comparison in the tabular domain remains limited. This understanding is crucial for both groups who use adversarial samples and may choose different attack strategies based on their goals, resources, and constraints; Defenders interested in assessing their model’s vulnerability often use adversarial attacks for automated penetration testing or adversarial training, which involves creating adversarial samples to improve the model’s robustness [19, 20]. Since defenders have full access to their model and can make unlimited queries, query-based attacks are a practical option.
In contrast, attackers aiming to minimize the number of queries performed and avoid detection, might prefer transferability-based attacks that use gradient methods on neural networks, as they capture underlying feature interactions which is necessary for subtle perturbations. However, systematic comparison of these attack strategies in the tabular domain remains absent.
To address these gaps and advance adversarial research in the tabular domain, a novel and comprehensive methodological framework that combines attack design and evaluation tailored specifically to tabular data characteristics was used.
To address the first gap of maintaining coherence in dependent features, this research develops a novel technique for consistent perturbation, using regression models to preserve realistic relationships in features with complex dependencies. This technique is designed for scenarios where these dependencies are not circular, ensuring that modifications maintain the inherent structure and correlations in the data.
To address the second gap of current evaluation methods’ limitations, the paper proposes two complementary evaluation criteria for assessing the quality of adversarial samples. The first criterion detects anomalies in feature space values, quantifying structural deviations from the data distribution. The second examines how adversarial samples influence the model’s decision-making process, employing SHapley Additive exPlanations (SHAP) [21], a widely adopted explainability (XAI) technique [22], to quantify inconsistencies in feature importance caused by adversarial perturbations. These measures address the limitations of relying solely on perturbation magnitude as an indicator of detectability.
Both criteria are employed using the proposed Class-Specific Anomaly Detection (CSAD) approach, which evaluates of adversarial samples relative to the benign distribution of their predicted class, rather than relying on global thresholds. This class-aware perspective improves sensitivity to inconsistencies in the tabular domain, where perturbations can be subtle and might be overlooked in global analyses.
To address the third gap of limited systematic comparison of attack strategies, a comprehensive evaluation of various black-box attack methods in the tabular domain is conducted. The evaluation integrates traditional metrics for computing attacker risk (success rate, query count, and $L_{p}$ norms) and attacker effort, complemented by the proposed quality criteria: sample anomaly detection and feature importance consistency. This thorough evaluation provides insights into both attack effectiveness and adversarial sample quality that can guide the development of more robust attacks and defenses.
The experiments performed in this research aim to answer the following research questions:
  • RQ#1 How do the perturbed samples of different attack strategies differ in terms of the attacker’s risk and effort?
  • RQ#2 To what extent do query-based attacks reflect the behavior of transferability-based gradient attacks?
  • RQ#3 For both benign samples and perturbed samples, is the attacked model’s decision-making process aligned with the expected feature importance?
The evaluation performed examines seven attack strategies across four tree-based classification models using three datasets: the Hateful Users on Twitter [23] and Intensive Care Unit [24] public datasets, which are benchmark datasets in the tabular domain, and a proprietary Video Transmission Quality dataset. The examined attack strategies consist of two query-based attacks (boundary and HopSkipJump [25, 26]) and five transferability-based gradient attacks with varying feature selection techniques [8, 9]. The process illustrated in  Fig.˜1.
@@
The contributions of this work can be summarized as follows:
  1. 1.
@@
    Proposing a novel technique for perturbing interdependent features in tabular data, leveraging regression models to preserve coherence and ensure feature consistency.
  2. 2.
@@
    Introducing CSAD, an approach that assesses anomalies relative to class-specific distributions rather than global thresholds, which improves sensitivity to feature inconsistencies.
  3. 3.
@@
    Developing comprehensive dual evaluation criteria based on anomaly assessment of the sample values and the feature importance, providing a principled framework for quantifying adversarial sample quality and its impact on a model’s decision-making process.
  4. 4.
@@
    Systematically comparing adversarial attack strategies’ performance factors in the tabular domain, providing critical insights into their effectiveness, limitations, and corresponding defense strategies. To the best of our knowledge, this is the first study to do this.
@@
The remainder of this paper is organized as follows: Section˜2 provides an overview of related works. Section˜3, which describes the methodological framework, is followed by Section˜4 and Section˜5, which present the experimental setup and provide a thorough analysis of the examined query- and transferability-based attack behavior. Section˜6 analyzes the trade-offs between different attack strategies, examines the implications for both attackers and defenders, and evaluates the effectiveness of detection mechanisms in tabular data domains. Finally, Section˜7 concludes the paper, providing key insights and future research directions for attackers and defenders in tabular data domains.
@@ 2 Background and Related Work
2.1 Adversarial Attacks in the Tabular Domain
@@
Szegedy et al. [27] introduced the concept of an adversarial attack in which an attacker provides a malicious data sample aimed at misleading the ML model and causing it to produce incorrect output. Their adversarial sample was crafted by performing subtle changes to an image, which were imperceptible to the human eye but misled the target model. However, research conducted in tabular subdomains, such as fraud detection and recommendation systems [6, 7, 28], has highlighted critical challenges within the tabular domain. These challenges include the constraints of editable and non-editable features (e.g., identification number, date of birth), data imbalance, and the presence of noncontinuous features (e.g., categorical features). The authors of these studies also mentioned the complexity of perturbing discrete values while maintaining the semantic integrity of the sample, as well as the challenge of capturing nuances when designing attacks. This discussion was further elaborated on in [8].
Most recent adversarial attacks operate under black-box conditions, where attackers lack prior knowledge of the target model’s architecture and parameters, possessing only query access [29]. Two main types of black-box attacks are query-based attacks and transferability-based attacks. In query-based attacks, including decision-based methods like the boundary Attack and HopSkipJump [25, 26], attackers iteratively adjust sample values based on model output until the optimization objective is fulfilled - misleading the target model to produce the desired output. These attacks typically begin with large initial modifications to the original sample, causing incorrect predictions, then optimize the modified sample to remain as close as possible to the original while maintaining the deception. In contrast, transferability-based attacks leverage surrogate models to craft an adversarial sample applying a white-box attack on the surrogate model, and then utilize the adversarial sample crafted to attack the target model in a single query [30].
Recent developments in tabular adversarial attacks have focused on addressing domain-specific constraints identified in earlier work. Cartella et al.  [7] addressed the limitation of applying image-based attacks to tabular data by modifying decision-based attacks to incorporate tabular data validity and editability constraints. Building on this foundation, Mathov et al.  [8] tackled the challenge of maintaining feature correlations, proposing an untargeted, transferability-based gradient attack that incorporats neural network (NN) embedding functions to preserve feature correlations and ensure value consistency by minimizing the $L_{2}$ distance and adjusting correlated features. Grolman et al. [9] extended this approach by integrating feature importance into the perturbation process, selecting the most important features for perturbation. Ju et al.  [31] further advanced the field by addressing feature distribution challenges, developing a transferability-based approach that considers feature distribution in the perturbation process, although validity constraints remained unaddressed.
2.2 Evaluation Challenges and Anomaly Detection
Current evaluation frameworks for adversarial attacks in various domains rely primarily on four dimensions: attack success rates [32, 6]; the number of queries made to the target model [33, 34]; computational efficiency, measured by the time required to create the adversarial sample, as well as the amount of data and time needed to train the surrogate model [35]; and perturbation magnitude, typically quantified through $L_{p}$ norms such as $L_{0}$, $L_{2}$, and $L_{\infty}$ [32, 36, 37, 6, 8, 9] or the average error between the original and adversarial samples [38].
In domains such as computer vision, perturbation magnitude is often used as an indicator of the imperceptibility of an adversarial sample. Specifically, an adversarial sample is considered indistinguishable from a benign sample if the human eye cannot perceive any difference from the original [27, 32, 39, 36]. In such cases, measuring perturbation magnitude through norms such as $L_{2}$ or $L_{\infty}$ is meaningful, because these norms capture relative changes in pixel values that directly influence visual detectability.
However, there are inherent limitations when employing perturbation magnitude as an evaluation measure in the tabular domain. While $L_{0}$, which counts the number of modified features, is particularly relevant, as it reflects the practical feasibility and potential cost of modifying specific attributes in real-world structured datasets, the applicability of $L_{2}$ and $L_{\infty}$ is restricted. These norms are only meaningful when all input features are numeric and share a common scale. More importantly, unlike images, where human visual inspection can directly assess perturbations, identifying such changes in tabular data requires domain expertise [6].
Existing works in the tabular domain have attempted to address this by involving domain experts who examine a small subset of “important” features, selected based on their significance to the model, to detect manipulations [6, 7]. However, this approach is neither objective, scalable, nor practical, as it requires manual assessment of each sample individually and does not account for the fact that altering even a small number of features can cause inconsistencies with the rest of the sample, making it distinguishable from benign data distributions.
Moreover, in the tabular domain, imperceptibility should not rely on the perturbation size but should instead critically depend on maintaining coherence and preserving the natural relationships and dependencies between features. Existing metrics fail to capture a sample’s coherence and feature consistency, both of which are essential for determining the true distinguishability of tabular adversarial samples.
To address these evaluation challenges, we propose using anomaly detection as a practical and scalable approach for the objective assessment of the distinguishability of adversarial samples in the tabular domain. The underlying assumption is that adversarial samples lacking structural coherence or consistency with benign data distributions are more likely to exhibit anomalous patterns, making them both distinguishable and detectable. However, the selection of suitable anomaly detection techniques for tabular data requires careful consideration of domain-specific constraints and data characteristics.